OpenStudy (anonymous):

find the tangents of the following curves x^2 + 16y^2 = 32 from (-6, 7/2). the answers are x+4y -8 and 41x + 4y + 232

OpenStudy (anonymous):

if you\[\frac{ dy }{ dx }\]what does that get?

OpenStudy (anonymous):

-x/6y

OpenStudy (anonymous):

\[\frac{ d x^2 }{ dx }=2x...\frac{d9}{dx}=0\] \[\frac{ dy }{ dx } = ??\]

OpenStudy (anonymous):

2x + 32y dy/dx = 0

OpenStudy (phi):

so you get \[ y'= \frac{-x}{16y} \]

OpenStudy (anonymous):

yes im sorry. typed it wrong.

OpenStudy (phi):

y' will be the slope of the tangent line at point (x,y) on the ellipse you want that slope to be the same as the slope between point (x,y) and point(-6,7/2) in other words: \[ \frac{y- \frac{7}{2}}{x+6}= \frac{-x}{16y} \]

OpenStudy (anonymous):

yes thats what i did and i got stuck up

OpenStudy (phi):

after cross multiplying you get \[ 16y^2 -56y = -x^2 -6x \\ x^2 + 16y^2 -56y +6x= 0 \] we can simplify that a bit... notice that the equation of the ellipse says \( x^2+16y^2=32\) so you have \[ 32 -56 y + 6x= 0 \\ 3x= 28y-16 \]

OpenStudy (phi):

remember that last equation, we will need it later. meanwhile, to continue, we must make that an equation of 1 variable. One way to proceed is square both sides (because we can know \( x^2 = 32-16y^2 \) from the equation of the ellipse squaring both sides: \[ 9x^2 = (28y-16)^2 \] replace x^2 to get \[ 9(32-16y^2) = (28y-16)^2 \]

OpenStudy (phi):

now solve for y.... use \( 3x= 28y-16\) to find x use x and y to find the slope using \( m= \frac{-x}{16y} \) now use the point - slope formula \[ y - y_0 = m(x - x_0) \] where \( (x_0,y_0) = ( -6,7/2) \) change to standard form to match your answers.

OpenStudy (anonymous):

thank you so much @phi !!